# DEGAP: Dual Event-Guided Adaptive Prefixes for Templated-Based Event Argument Extraction Model with Slot Querying

### v1

@@ 3.1.4 Encoding with Prefixes
To ensure a fair comparison and to visually demonstrate the effects of our designed prefixes, we adopt the encoding process from prior work He et al. (2023) for encoding context instance and event template, and augment it with our prefixes.
Given input context $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and the event trigger $x_{trig}$ triggering event type $e$, we use text markers <t> and </t> as special tokens and then insert them before and after $x_{trig}$ in $X$, respectively.
@@
Then, we input the processed $\widetilde{X}$ into the encoder and the decoder to obtain the context representation $H_{x}$:
$H_{x}^{enc}=\text{Encoder}(\text{Prefix}_{enc}^{ins},\widetilde{X}),$ (6)
$H_{x}=\text{Decoder}(\text{Prefix}_{dec}^{ins},H_{x}^{enc}),$ (7)
@@
$H_{\mathcal{T}_{e}}=\text{Decoder}(\text{Prefix}_{dec}^{tem},H_{\mathcal{T}_{e}}^{enc},H_{x}^{enc}),$ (9)
where $\text{Prefix}_{enc}^{tem}$ and $\text{Prefix}_{dec}^{tem}$ are the template-oriented prefix representations $\text{Prefix}_{i}^{tem}$ from the encoder and decoder, respectively. Note that the cross-attention mechanism is only utilized to decode the event template. When computing Eq. 7, it is skipped.
@@ 3.1.5 Span Selection
$\text{logit}_{k}^{start}=\phi_{k}^{start}H_{x}\in\mathbb{R}^{n},$
$\text{logit}_{k}^{end}=\phi_{k}^{end}H_{x}\in\mathbb{R}^{n},$
@@
where candidate span set $\mathcal{C}=\{(l,r)\mid(l,r)\in n^{2},0<r-l<MSL\}\cup\{(0,0)\}$. $MSL$ denotes the max span length.
The loss function is defined as:
@@
where $X$ ranges over all context in dataset and $k$ ranges over all role slots in the template for $X$, and $(s_{k},e_{k})$ is the golden span for role slot $k$. For a role slot that does not relate to any argument, it is assigned the empty span $(0,0)$.
@@ 4.1 Experimental Setup
Model PLM ACE05 RAMS WIKIEVENTS MLEE
@@
TSAR Xu et al. (2022) RoBERTa - - 57.0 52.1 71.1 65.8 72.6 71.5
SCPGR Liu et al. (2023c) RoBERTa - - 56.5 52.2 71.5 66.2 72.5 71.3
BART-Gen Li et al. (2021) RoBERTa 69.9 66.7 51.2 47.1 66.8 62.4 71.0 69.8
@@
TabEAE He et al. (2023) RoBERTa 75.5 72.6 57.0 52.5 70.8 65.4 71.9 71.0
DRAAE (Ours) RoBERTa 76.1 73.2 57.5 53.0 71.0 66.0 72.8 71.7
DEGAP (Ours) RoBERTa 76.6 74.4 58.5 54.2 72.2 67.1 74.0 73.4
@@
Datasets. We conduct experiments on four benchmark datasets, specifically including ACE05 Doddington et al. (2004), RAMS Ebner et al. (2020), WIKIEVENTS Li et al. (2021), and MLEE Pyysalo et al. (2012). ACE05 is a sentence-level dataset, while the other three are document-level datasets. We provide detailed introductions to these datasets in Appendix A.
@@
Implementation details. Please refer to Appendix B.1 for details of implementation.
Baselines. Details about the baselines are listed in Appendix B.2. For a comprehensive comparison, we also conduct a comparative analysis between DEGAP and LLMs in Appendix C.1.
We also compare our method with its variant, the Dual Retrieval-Augmented Argument Extraction model (DRAAE), which retrieves relevant event instances and templates as cues to guide the model, instead of using the prefixes we designed. We provide further implementation details of this model in Appendix B.3.
@@ 4.2 Overall Performance
Table 1 presents the performance of all baselines and our model across all datasets. From the results, we can conclude:
@@
(1) Our method significantly outperforms all the baselines and achieves new SOTA results on the four datasets. Specifically, our method improves the Arg-C F1 score by 1.0%, 0.9%, 0.9%, and 1.9% on ACE05, RAMS, WIKIEVENTS, and MLEE, respectively. This demonstrates the efficacy of our method in handling both sentence-level EAE and document-level EAE tasks.
(2) Compared to methods using retrieval (i.e., R-GQA and AHR) and those using event templates (e.g., DEGREE and TabEAE), our approach shows better performance on all datasets. We attribute this to DEGAP’s ability to fully utilize relevant event instance and event template information as cues. In contrast, previous methods provide insufficient guidance to the model from only one perspective.
@@
(3) Compared to methods that use prefixes, such as AMPERE and SPEAE, which utilize auxiliary semantic information to augment their prefixes, our method does not rely on these modules and still surpasses them to varying degrees. This suggests that our event-guided prefixes are more suitable for EAE, thereby leveraging their advantages.
(4) Our DRAAE surpasses the baseline model TabEAE across all datasets. This indicates that augmenting input with relevant event instances and event templates is effective for extracting arguments. However, our DEGAP still outperforms DRAAE, primarily because it addresses two shortcomings in retrieval: a) the need to pre-specify the number of relevant instances and templates to be retrieved, and potentially irrelevant retrieval results, leading to insufficient utilization of relevant semantic information; b) the utilization of pre-trained retriever with frozen parameters, leading to less relevance to the EAE task and yielding a poor local optimum. Moreover, the efficiency analysis in Appendix C.2 indicates that due to our DEGAP not requiring retrieval, it has a faster inference speed.
@@
5 Analysis
5.1 Why is DEGAP Effective?
We explore from two perspectives: a) domain transfer and b) handling complex contextual semantics, to investigate whether the performance gains in our method are due to DEGAP providing the model with relevant event instances and event templates as cues. We consider the base model TabEAE.
Domain Transfer. We examine the performance of the model when facing domain transfer. Specifically, we select four different domain datasets: ACE05, RAMS, WIKIEVENTS, and MLEE, train the model on one domain dataset (src), and then test it on all domain datasets (tgt). The results are shown in Table 2. Compared to the results w/o domain transfer, both our method DEGAP and the baseline TabEAE exhibit a sharp decline under different src-tgt transfer settings, where the baseline’s performance even falls below 10% in some cases (e.g., WIKI $\Rightarrow$ MLEE). This indicates that extracting arguments from new types of events is highly challenging for the model. However, encouragingly, our method DEGAP consistently outperforms the baseline in different src-tgt transfer settings. We attribute this to DEGAP’s ability to use relevant semantic information in the source domain as cues and learn by analogy to better understand new types of events in the target domain.
@@ 5.1 Why is DEGAP Effective?
Handing Complex Contextual Semantics. We explore the ability of models to handle complex semantics from two perspectives: the number of events and the degree of event relevance. (1) Number of events. Multiple events in the same context often exhibit complex semantic connections. We measured the performance of the models on instances whose context contains different numbers of events. As shown in Table 3, compared to the baseline TabEAE, our method shows significant improvements in performance across all datasets, especially in instances that contain multiple events within the context. (2) Degree of event relevance. If two events in the context share arguments, their semantic connection is often more complex. We measured the performance of the model in extracting arguments from overlapping events (events with shared arguments). Table 4 shows that our method achieves the best performance across all datasets, particularly gaining greater improvements in handling overlapping events.
Experimental results support our conclusion: by fully utilizing relevant event instances and event templates to guide the model, our method excels in identifying the target event from complex contextual semantics.
5.2 Why are Dual Prefixes Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 DEGAP-CCT 73.6 53.7 66.6 72.3
3 DEGAP-SP 73.9 53.4 66.8 72.6
4 only IOP (w/o TOP) 73.8 52.9 66.4 72.6
5 only TOP( w/o IOP) 73.6 52.7 66.3 72.7
6 w/o DEGAP 72.8 52.0 65.7 71.2
7 DEGAP-TST 73.9 53.0 66.6 72.8
  • DEGAP-CCT (Concatenate Context and Template): Instead of encoding the context and template separately, this variant directly concatenates them and then inputs the combined sequence into the Encoder, which is equipped with a set of shared prefixes. For a fair comparison, we use RoBERTa-large as the encoder.
  • DEGAP-SP (Shared Prefix): This variant follows the same encoding process as DEGAP but replaces the instance-oriented and template-oriented prefixes with a single set of shared prefixes.
Moving forward, we further investigate the effects of instance-oriented prefix (IOP) and template-oriented prefix (TOP). Finally, we consider DEGAP-TST (event-Template-Specific TOP), where each event template has its specific TOP. Therefore, TOP solely learns and provides semantic information specific to the target event template, without considering other relevant event templates. Table 5 shows the results.
From the table, we observe that compared to our DEGAP (line 1), the single-prefix variants DEGAP-CCT and DEGAP-SP (lines 2-3) show performance declines to varying degrees across all datasets. We speculate that this may be due to the prefixes becoming confused by simultaneously learning information about both event templates and event instances. Therefore, we choose to use dual prefixes instead of single prefixes in the proposed method.
Additionally, compared to the baseline w/o DEGAP (line 6), introducing IOP or TOP separately will result in a performance increase on all datasets (lines 4-5), and the performance gains by introducing them at the same time exceed the performance improvement caused by introducing any one module alone (line 1). This indicates that it is necessary to utilize IOP and TOP to respectively provide the model with relevant event instance information and event template information as cues for inference. Moreover, IOP and TOP can offer complementary guidance to the model from the event mention-level and ontology level, respectively.
Finally, an interesting phenomenon is that despite DEGAP-TST introducing more parameters than DEGAP, its performance declines across all four datasets (line 7). This further confirms the necessity of providing the model with relevant event template information as cues.
5.3 Is Event-Guided Adaptive Gating mechanism Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 w/o EGAG 73.4 53.3 66.4 72.3
3 DEGAP-S 73.9 53.6 66.8 72.6
To explore the importance of the Event-Guided Gating Adaptive mechanism (EGAG), we compare DEGAP with the following variants:
  • w/o EGAG: To investigate whether the performance improvement of DEGAP is due to the additional parameters introduced (gating weights and scaling factors) or because the event-guided prefixes are truly more suitable for EAE, we designed the variant DEGAP w/o EGAG. For a fair comparison, we expanded the prefix length to 1.5 times to align its parameter count with DEGAP.
  • DEGAP-S: To validate the importance of using the target event’s trigger and type to guide IOP and TOP respectively, we follow Zhang et al. (2023d) to directly use the hidden state of the first input token “<s>” in the previous layer as $h_{i-1}^{ins}$ in Eq. 1 and $h_{i-1}^{tem}$ in Eq. 4, respectively.
Table 6 shows that removing EGAG results in a noticeable decline in Arg-C F1 scores across the four datasets. This phenomenon indicates that event-guided prefixes are indeed more suitable for EAE. Additionally, DEGAP-S also shows a decrease in Arg-C F1 scores across the four datasets. This may be because it uses only the first input token “<s>” to calculate $h_{i-1}^{ins}$ and $h_{i-1}^{tem}$, which is insufficient to capture information relevant to the target event in the prefixes. However, compared to the model w/o EGAG, DEGAP-S shows improved performance across all four datasets. This further confirms that previous prefixes do not fully meet the specific informational needs of EAE and leverage their advantages.
We further demonstrate the effects of EGAG visually on the ACE05 dataset.
Attention Visualization. We randomly select a sample and visualize the weight distribution of prefix tokens in IOP and TOP. As shown in Fig. LABEL:prefix_token_weight, the target event has different levels of attention to the prefixes of different layers. More specifically, even within the same layer, the attention to different prefix tokens varies. This observation confirms that EGAG can adaptively capture information from the prefix tokens at each layer based on the target event and further guide the prefixes.
Prefix length. We further analyze the impact of prefix length on model performance. Intuitively, a longer prefix implies more trainable parameters, thus enabling it to provide more expressive guidance for the model regarding relevant semantic information. As shown in Fig. LABEL:prefix_len, as the prefix length increases, the model performance tends to improve, followed by a slight decline. We speculate that this is because overly long prefixes provide redundant semantic information, making it difficult for the model to focus on extracting the target event.
6 Conclusion
In this work, we introduce DEGAP. Based on observations of existing research and comprehensive experimental analysis, we show that prefixes can effectively provide relevant event information to models as cues. Additionally, we introduce an event-guided adaptive gating mechanism that guides prefixes to adapt to different events to fully leverage their advantages. Through extensive experiments, we demonstrate the validity of our newly introduced methods. In the future, we plan to integrate more beneficial auxiliary information into prefixes for solving EAE.
Limitations
We summarize the limitations of our method as follows:
  • Although experimental results indicate that our method exhibits profound performance, our designed prefixes lack sufficient interpretability compared to manually crafted prompts.
  • Due to limited computational resources, we only equipped our prefixes for use on PLMs. In fact, some other works have successfully integrated these efficient prefix parameters into LLMs (e.g., Llama 2 Touvron et al. (2023)) and demonstrated excellent results Zhang et al. (2023c). How to effectively combine our prefixes with these LLMs for solving the EAE task remains an open research question.
  • Compared to other methods, our approach requires additional time costs to find suitable prefix lengths for both IOP and TOP.
  • While our approach is immune to high computational costs and inefficient use of information from relevant instances and templates due to not requiring retrieval, it also presents a significant challenge in how to benefit from other useful semantic information beyond the training corpus. In contrast, retrieval-based methods can leverage richer external semantic information by expanding the retrieval source.
We will attempt to address these issues in future work.
Appendix A Datasets

### v2

@@ 3.1.4 Encoding with Prefixes
To ensure a fair comparison and to visually demonstrate the effects of our prefixes, we adopt the encoding process from prior work He et al. (2023) for encoding context instance and event template, and augment it with our prefixes. With both relevant event instances and templates as cues , model can better understand the target event from context.
Given input context $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and the event trigger $x_{trig}$ , we use text markers <t> and </t> as special tokens and then insert them before and after $x_{trig}$ in $X$, respectively.
@@
Then, we input the processed $\widetilde{X}$ into the encoder and the decoder to obtain the context representation $H_{\widetilde{X}}$:
$H_{\widetilde{X}}^{enc}=\text{Encoder}(\text{Prefix}_{enc}^{ins},\widetilde{X}),$ (6)
$H_{\widetilde{X}}=\text{Decoder}(\text{Prefix}_{dec}^{ins},H_{\widetilde{X}}^{enc}),$ (7)
@@
$H_{\mathcal{T}_{e}}=\text{Decoder}(\text{Prefix}_{dec}^{tem},H_{\mathcal{T}_{e}}^{enc},H_{\widetilde{X}}^{enc}),$ (9)
where $\text{Prefix}_{enc}^{tem}$ and $\text{Prefix}_{dec}^{tem}$ are the template-oriented prefix representations $\text{Prefix}_{i}^{tem}$ from the encoder and decoder, respectively.
@@ 3.1.5 Span Selection
$\text{logit}_{k}^{start}=\phi_{k}^{start}H_{\widetilde{X}}\in\mathbb{R}^{n},$
$\text{logit}_{k}^{end}=\phi_{k}^{end}H_{\widetilde{X}}\in\mathbb{R}^{n},$
@@
where candidate span set $\mathcal{C}=\{(l,r)\mid(l,r)\in n^{2},0<r-l<MSL\}\cup\{(0,0)\}$. $MSL$ denotes the max span length. The loss function is defined as:
@@
where $X$ ranges over all contexts in dataset and $k$ ranges over all role slots in the template for $X$, and $(s_{k},e_{k})$ is the golden span for role slot $k$. For a role slot that does not relate to any argument, it is assigned the empty span $(0,0)$.
@@ 4.1 Experimental Setup
Category Model PLM ACE05 RAMS WIKIEVENTS MLEE
@@
Retrieval-based R-GQA Du and Ji (2022) BART 75.6 72.8 - - - - - -
AHR Ren et al. (2023) T5 - - 54.6 48.4 69.6 63.4 - -
Templated-based BART-Gen Li et al. (2021) RoBERTa 69.9 66.7 51.2 47.1 66.8 62.4 71.0 69.8
@@
PAIE Ma et al. (2022) BART 75.7 72.7 56.8 52.2 70.5 65.3 72.1 70.8
TabEAE He et al. (2023) RoBERTa 75.5 72.6 57.0 52.5 70.8 65.4 71.9 71.0
Prefix-based APE Zhang et al. (2023b) BART 75.3 72.9 56.3 51.7 70.6 65.8 - -
@@
Retrieval-based DRAAE (Ours) RoBERTa 76.1 73.2 57.5 53.0 71.0 66.0 72.8 71.7
Prefix-based DEGAP (Ours) RoBERTa 76.6 74.4 58.5 54.2 72.2 67.1 74.0 73.4
@@
Datasets. We conduct experiments on four benchmark datasets, specifically including ACE05 Doddington et al. (2004), RAMS Ebner et al. (2020), WIKIEVENTS Li et al. (2021), and MLEE Pyysalo et al. (2012). ACE05 is a sentence-level dataset, while the other three are document-level datasets. We provide detailed introductions to these datasets in Appendix C.1.
@@
Implementation details. Please refer to Appendix C.2 for details of implementation.
Baselines. Details about the baselines are listed in Appendix C.3. For a comprehensive comparison, we also conduct a comparative analysis between DEGAP and LLMs in Appendix D.1, showing that DEGAP can outperform the latest LLMs for EAE.
We also compare our method with its variant, the Dual Retrieval-Augmented Argument Extraction model (DRAAE), which only retrieves relevant event instances and templates as cues to guide the model, instead of using the prefixes we designed. We provide further implementation details of this model in Appendix C.4.
@@ 4.2 Overall Performance
Table 1 presents the performance of all baselines and our method across all datasets. Our method outperforms all the baselines and achieves new SOTA results on the four datasets. Specifically, our method improves the Arg-C F1 score by 1.0%, 0.9%, 0.9%, and 1.9% on ACE05, RAMS, WIKIEVENTS, and MLEE, respectively. This demonstrates the efficacy of our method in handling both sentence-level EAE and document-level EAE tasks. Moreover, we can conclude:
(1) Augmenting input with both relevant event instances and templates is beneficial for extracting arguments. Our DRAAE and DEGAP outperform previous retrieval-based and template-based methods, which only utilize event instances or templates to provide insufficient guidance for model. Our approach can fully combine these beneficial cues to better understand the target event from context.
@@
(2) Our event-guided prefix can effectively coping with the specific informational needs of different events, which is superior to previous prefixes. Compared to previous prefix-based methods, such as AMPERE and SPEAE, which utilize auxiliary semantic information to augment their task-general prefixes, our method does not rely on these modules and still surpasses them to varying degrees. This reveals that by using EGAG to connect other relevant events to the target event, our method can effectively utilize relevant information in the prefix and discard irrelevant information.
(3) Our prefixes can utilize relevant event instances and templates more sufficiently and efficiently to boost performance compared to retrieval. As shown in Table 1, Our DEGAP outperforms DRAAE on all datasets. This is because 1) DRAAE uses a pre-trained retriever with frozen parameters to retrieve event instances related to the target event, but fails to find the most relevant events, and 2) DRAAE retrieves and encodes independently, leading to a poor local optimum for each of the steps. In contrast, our prefixes are trained and updated along with model parameters and can fully learn the semantic information of different event instances and templates, avoiding these issues. Moreover, the efficiency analysis in Appendix D.2 indicates that DEGAP has a faster inference speed due to the removal of retrieval.
5 Analysis
5.1 Why is DEGAP Effective?
We explore from two perspectives: 1) domain transfer and 2) handling complex contextual semantics, to investigate whether the performance gains in our method are due to DEGAP providing the model with relevant event instances and event templates as cues. We consider the base model TabEAE, which primarily differs from our model in whether dual prefixes are used.
Domain Transfer. We assume that DEGAP can use relevant event information from the source domain as cues and learn by analogy to better understand new types of events in the target domain. To this end, we examine the model performance when facing domain transfer. Specifically, we select four different domain datasets: ACE05, RAMS, WIKIEVENTS, and MLEE, train the model on one domain dataset (src), and then test it on all domain datasets (tgt). The results are shown in Table 4 in Appendix A.1. Compared to the results w/o domain transfer, both our method DEGAP and the baseline TabEAE exhibit a sharp decline under different src-tgt transfer settings, where the baseline’s performance even falls below 10% in some cases (e.g., WIKI $\Rightarrow$ MLEE). This indicates that extracting arguments from new types of events is highly challenging for the model. However, encouragingly, our method DEGAP consistently outperforms the baseline in different src-tgt transfer settings. This fully confirms that our hypothesis is correct.
Handing Complex Contextual Semantics. We believe that by fully utilizing relevant event instances and event templates to guide the model, our method excels in identifying the target event from complex contextual semantics. To this end, we explore the ability of models to handle complex semantics from two perspectives: the number of events and the degree of event relevance. (1) Number of events. Multiple events in the same context often exhibit complex semantic connections. We measured the performance of the models on instances whose context contains different numbers of events. As shown in Table 5 in Appendix A.2, compared to the baseline TabEAE, our method shows consistent improvements in performance across all datasets, especially in instances that contain multiple events within the same context. (2) Degree of event relevance. If two events in the context share arguments, their semantic connection is often more complex. We measured the performance of the model in extracting arguments from overlapping events (events with shared arguments). Table 6 shows that our method achieves the best performance across all datasets, particularly gaining greater improvements in handling overlapping events. Experimental results support our conclusion
5.2 Why are Dual Prefixes Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 DEGAP-CCT 73.6 53.7 66.6 72.3
3 DEGAP-SP 73.9 53.4 66.8 72.6
4 only IOP (w/o TOP) 73.8 52.9 66.4 72.6
5 only TOP ( w/o IOP) 73.6 52.7 66.3 72.7
6 w/o DEGAP 72.8 52.0 65.7 71.2
7 DEGAP-TST 73.9 53.0 66.6 72.8
(1) DEGAP-CCT (Concatenate Context and Template): Instead of encoding the context and template separately, this variant directly concatenates them and then inputs the combined sequence into the encoder, which is equipped with a shared prefix. For a fair comparison, we use RoBERTa-large as the encoder.
(2) DEGAP-SP (Shared Prefix): This variant follows the same encoding process as DEGAP but replaces instance-oriented and template-oriented prefixes with a shared prefix.
Moving forward, we further investigate the effects of instance-oriented prefix (IOP) and template-oriented prefix (TOP). Finally, we consider DEGAP-TST (event-Template-Specific TOP), where each event template has its specific TOP. Therefore, each TOP solely learns and provides semantic information specific to the target event template, leading to insufficient interaction between event templates. Table 2 shows the results.
From the table, we observe that compared to our DEGAP (line 1), the single-prefix variants DEGAP-CCT and DEGAP-SP (lines 2-3) show performance declines to varying degrees across all datasets. This may be because the prefix simultaneously learns the semantic information of event instances and templates, leading to confusion between this information and failing to accurately provide sufficient guidance for model when encoding context and event template. In contrast, dual prefixes separately learn semantic information of event instance and template, which can better provide useful cues for model when encoding context and event template.
Additionally, compared to the baseline w/o DEGAP (line 6), introducing IOP or TOP separately will result in a performance increase on all datasets (lines 4-5), and the performance gains by introducing them at the same time exceed the performance improvement by introducing any one module alone (line 1). This indicates that it is necessary to utilize IOP and TOP to respectively provide the model with relevant event instance information and event template information as cues for inference. Moreover, IOP and TOP can offer complementary guidance to the model from the event mention-level and ontology level, respectively.
Finally, an interesting phenomenon is that despite DEGAP-TST introducing more parameters than DEGAP, its performance declines across all datasets (line 7). This confirms the importance of considering the potential connections between templates, where some relevant templates can usually provide useful information to the target event.
5.3 Is Event-Guided Adaptive Gating mechanism Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 w/o EGAG 73.4 53.3 66.4 72.3
3 DEGAP-S 73.9 53.6 66.8 72.6
To explore the importance of the Event-Guided Adaptive Gating mechanism (EGAG), we compare DEGAP with the following variants:
(1) w/o EGAG: To investigate whether the performance improvement of DEGAP is due to the additional parameters introduced (gating weights and scaling factors) or because event-guided prefixes can better provide model with relevant event information as cues, we designed the variant DEGAP w/o EGAG. For a fair comparison, we expanded the prefix length to 1.5 times to align its parameter count with DEGAP.
(2) DEGAP-S: To validate the importance of using the target event to guide prefix, we follow Zhang et al. (2023d) to directly use the sentence representation in the previous layer as $h_{i-1}^{trig}$ in Eq. 1 and $h_{i-1}^{type}$ in Eq. 3, respectively.
Table 3 shows that removing EGAG results in a noticeable decline in Arg-C F1 scores across the four datasets. This phenomenon indicates that utilizing EGAG to connect other relevant events to the target one is very necessary, as it allows our method to further effectively utilize relevant information in the prefix and discard irrelevant information. Additionally, DEGAP-S also shows a decrease in Arg-C F1 scores across the four datasets. This may be because it uses only the first input token “<s>” to calculate $h_{i-1}^{trig}$ and $h_{i-1}^{type}$, which is insufficient to connect other relevant events to the target event and thus capture relevant information from the prefix. However, compared to the model w/o EGAG, DEGAP-S shows improved performance across all four datasets. This further confirms the importance of utilizing EGAG to connect other relevant events to the target event.
Attention Visualization. We further demonstrate the effects of EGAG visually on the ACE05 dataset. We randomly select a sample and visualize the weight distribution of prefix tokens in IOP and TOP. As shown in Fig. LABEL:prefix_token_weight in Appendix B, the target event has different levels of attention to the prefixes of different layers. More specifically, even within the same layer, the attention to different prefix tokens varies. This observation confirms that EGAG can adaptively capture the relevance of other event to the given one and further extract relevant information from the prefix tokens at each layer.
Prefix length. Moreover, We analyze the impact of prefix length on model performance in Appendix D.3.
6 Conclusion
In this work, we introduce DEGAP. Based on observations of existing research and comprehensive experimental analysis, we show that prefix can effectively learn semantic information from different event instances and templates, and allow them to interact fully. Additionally, we introduce an event-guided adaptive gating mechanism to connect other relevant events to the target event and thus capture relevant information from the prefix. Finally, these event-guided prefixes provide relevant event instances and template information as cues for the target event. Extensive experiments demonstrate the validity of our newly introduced methods. In the future, we plan to integrate other useful auxiliary information into prefixes for solving EAE.
Limitations
We summarize the limitations of our method as follows:
(1) Although experimental results indicate that our method exhibits profound performance, our designed prefixes lack sufficient interpretability compared to methods that directly retrieve relevant event instances and templates to enhance performance (e.g., DEGAP vs DRAAE). Additionally, compared to other methods, our approach requires additional time costs to find suitable prefix lengths for both IOP and TOP.
(2) Due to limited computational resources, we only equipped our prefixes for use on PLMs. In fact, some other works have successfully integrated these efficient prefix parameters into LLMs (e.g., Llama 2 Touvron et al. (2023)) and demonstrated excellent results Zhang et al. (2023c). How to effectively combine our prefixes with these LLMs for solving the EAE task remains an open research question.
We will attempt to address these issues in future work.
Appendix A Why is DEGAP Effective?
@@
A.1 Domain Transfer.
[…]
A.2 Handing Complex Contextual Semantics.
@@
We explore the ability of models to handle complex semantics from two perspectives: the number of events and the degree of event relevance. Table 5 presents the Arg-C F1 score on instances with different numbers of events in context, and Table 6 shows the results of extracting overlapping (events with shared arguments) and non-overlapping events. We can infer that the performance improvement of DEGAP is more attributed to event instances that contain multiple events within the same context and overlapping events. This may be because by fully utilizing relevant event instances and event templates to guide the model, our DEGAP can better identify the target event from complex contextual semantics.
Appendix B Attention Visualization
We demonstrate the effects of EGAG visually on the ACE05 dataset. Fig. LABEL:prefix_token_weight shows the weight distribution of prefix tokens guided by the target event. We can observe that the target event has different attentions on different layers and different prefix tokens within the same layer. This indicates that with EGAG, the target event is adaptively capturing its relationship with other events, thus further extracting relevant information from the prefix tokens at each layer.
Appendix C Experimental Setup
C.1 Datasets